Three charges are arranged as shown in the 1 igure. The value of Coulombs constant is 8.98755 x 10 N m2/C2 2.6 nC -0.8 nC 3.5 m 1.6 m ● 0.51 nC Find the magnitude of the electrostatic force on the charge at the origin. Answer in units of nN
Answer #1

Given,

q1 = -0.8 nC ; q2 = 2.6 nC ; q3 = 0.51 nC

r1 = 3.5 m ; r2 = -1.6 m

The electroststic force between q1 and q2 is along x

Fx = k q1q2/r1^2

Fx = 9 x 10^9 x 0.8 x 10^-9 x 2.6 x 10^-9/3.5^2 = 1.53 x 10^-9 N

that between 1 and 3 is along -y

Fy = 9 x 10^9 x 0.8 x 10^-9 x 0.51 x 10^-9/1.6^2 = 1.43 x 10^-9 N

The net force will be:

F = sqrt [(1.53 x 10^-9^2) + (1.43 x 10^-9)^2] = 2.1 x 10^-9 N = 2.1 nano Newtons

Hence, F = 2.1 nN